OpenStudy (anonymous):

Find a pair of factors for each number by using the difference of two squares. a. 45 b. 77 c. 112

OpenStudy (anonymous):

Fan and Medal Please? 45=49-4=(7+2)(7-2)= 9 x 5 77= 81-4= (9+2)(9-2)= 11x 7 112= 121-9= (11+3)(11-3)= 14 x 8

OpenStudy (anonymous):

I will later,,, the site keeps having to reconnect -.- and thank you